OpenStudy (anonymous):

Find second derivative of the function r={1+4theta / 4theta} (4-theta)

OpenStudy (anonymous):

\[r = \frac{1 + 4\theta}{4\theta} \times(4-\theta)\] ?

OpenStudy (anonymous):

yes, how to do second derivative please

OpenStudy (anonymous):

sure, you just take the derivative, then take it again on the derivative. you know how to thake the derivative of this fucntion r?

OpenStudy (anonymous):

No, I'm confused when it's quotient function, can you help

OpenStudy (anonymous):

sure, the easiest thing to do here would be to multiply by the 4 - theta first

OpenStudy (anonymous):

do that and type what you get in the equation editor

OpenStudy (anonymous):

simplify it as much as possible

OpenStudy (anonymous):

can you show it here so I can see?

OpenStudy (anonymous):

\[\frac{d}{dx}( \frac{g(x)}{f(x)}) = \frac{g(x)'f(x) - f(x)g(x)'}{f(x)^2}\]

OpenStudy (anonymous):

so what is g(x) for you?

OpenStudy (anonymous):

1+4theta rigt?

OpenStudy (anonymous):

no, simplify the equation first

OpenStudy (anonymous):

multiply 1 +4theta by (4 - theta)

OpenStudy (anonymous):

then you will just have r = somthing/something

OpenStudy (anonymous):

differentiate using the rule i gave you up there^

OpenStudy (anonymous):

oh ic, much better

OpenStudy (anonymous):

another question??

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

give me few minutes as I'm trying to solve for this problem, thanks

OpenStudy (anonymous):

well im going to go now. Good luck!

OpenStudy (anonymous):

thanks for your help